Monday, October 2, 2017

Solid progress on the HP 1000F system


Now that I am comfortable the processor power supply is working, I could put the racks together and mount the processor. This involved some power cabling to bring power to the slave cabinet, plus routing and installation of many cables into the processor.

IO cables installed onto processor boards
The floating point unit is contained in a separate box mounted just below the main processor. A ribbon cable snakes up and is connected to the back of the display panel. I had to guess about orientation since the cable and connector are not polarized.

Cable from floating point unit attached at front left

Cable from floating point routed upward to processor box
Most of the peripheral cables were obvious - terminals connect to the async IO boards, the HP-IB cable hooks to the bus controller card, mag tape connectors go to their board, the disk controller cable hooks to its card - but there is a mystery cable which 'might' connect to the "microcircuit" board and a RS-232 cable that doesn't seem to match any board's edge connector.

I powered up and the processor appeared to complete its power-on diagnostics successfully. I am waiting for more documentation but the advice from Jay West who has restored some of these systems on his own is that the system seems to have passed its self-test diagnostics. At this point, my working assumption is that the processor is healthy.

Power fail light warns that memory contents were not preserved
Hand toggled code can be executed as long as I can bear the tedium of this slow entry method.
The main power button on the rack began to stick in the off position, something I will need to fix to continue using the system. The switch itself has a lamp inside to indicate when the rack is energized. It is pushed in to turn off and pushed again to have it pop out to the 'on' position.

Sticky power switch for the rack enclosure
I 'hotwired' the rack around the switch while I work on it to make it work better. The tape drive and the disk controller also came up seemingly fine. I am not going to power on the 7906 disk drive until I have replaced any crumbling foam filters, cleaned the platters and heads carefully, and inspected everything. It is the only box that is not yet installed in the racks. 
HP 1000 F system, powered up and waiting for further testing

No comments:

Post a Comment